# Flexible Tails for Normalizing Flows

### v1

@@ Abstract
Normalizing flows are a flexible class of probability distributions, expressed as transformations of a simple base distribution. A limitation of standard normalizing flows is representing distributions with heavy tails, which arise in applications to both density estimation and variational inference. A popular current solution to this problem is to use a heavy tailed base distribution. Examples include the tail adaptive flow (TAF) methods of Laszkiewicz et al. (2022). We argue this can lead to poor performance due to the difficulty of optimising neural networks, such as normalizing flows, under heavy tailed input. This problem is demonstrated in our paper. We propose an alternative: use a Gaussian base distribution and a final transformation layer which can produce heavy tails. We call this approach tail transform flow (TTF). Experimental results show this approach outperforms current methods, especially when the target distribution has large dimension or tail weight.
@@ 1 Introduction
Modelling distributions with heavy tails is critical in many applications such as climate (Zscheischler et al., 2018), contagious diseases (Cirillo and Taleb, 2020) and finance (Gilli and Këllezi, 2006). Indeed, in the context of density estimation, results from extreme value theory (EVT) (Coles, 2001; Embrechts et al., 2013) show distributional tails can often be modelled using a particular heavy tailed distribution, the generalized Pareto distribution (GPD). Heavy tailed targets can also arise naturally in Bayesian inference, requiring heavy tailed approximate distributions in variational inference (Liang et al., 2022) (and providing challenges for MCMC methods—see e.g. Yang et al., 2022).
However standard normalizing flows do not model heavy tails well. In particular, Jaini et al. (2020) prove that Gaussian tails cannot be mapped to GPD tails under Lipschitz transformations. (This is a special case of their main result, which we present as Theorem 2 below.) Many normalizing flows use Lipschitz transformations of Gaussian base distributions, implying they produce distributions which are poor approximations of heavy tailed distributions.
Based on this result, several authors have proposed using heavy tailed base distributions. This includes the tail adaptive flow (TAF) methods of Laszkiewicz et al. (2022) for density estimation, and similar methods for variational inference (Liang et al., 2022). We argue that using heavy tailed base distributions has an important drawback. Typical NF transformations involve passing the input $z$ through a neural network. However neural network optimisation can perform poorly under heavy tailed input—as this can induce heavy tailed gradients, which are known to be problematic (Zhang et al., 2020). We verify this problem empirically, first in Section 4.1 for a simple neural network regression example, and then in Section 4.2 for normalizing flows.
To improve performance we propose an alternative approach: use a Gaussian base distribution with a final non-Lipschitz transformation in $T$. We refer to this as the tail transform flow (TTF) approach. The difference is represented schematically in Figure 1. In Section 3.1 we provide a suitable final transformation (3), which is motivated by extreme value theory. This is easy to implement in automatic differentiation libraries as it’s based on a standard special function: the complementary error function. We also prove that this transformation converts Gaussian tails to heavy tails with tunable tail weights, and hence provides the capability of producing heavy tails when used with standard flows.
Figure 1: Our proposed method changes the location where extremes/tails are modelled.
We perform experiments which show our TTF approach outperforms current methods, especially when the target distribution has large dimension or tail weight. We concentrate on density estimation, investigating both synthetic and real data. A proof-of-concept variational inference experiment is also included.
@@
  • We illustrate the problem of using extreme inputs in (a) generic neural networks (b) normalizing flows.
  • We introduce the TTF transformation—equation (3) below—and prove results to show it converts Gaussian tails to heavy tails with tunable tail weights. This provides the capability of producing heavy tails when used with standard NFs.
  • We provide practical methodology to use TTF in a normalizing flow architecture.
  • We demonstrate improved empirical results for density estimation (synthetic and real data examples) and variational inference (a proof-of-concept example with an artificial target) compared to standard normalizing flows, and other NF methods for heavy tails.
@@ 1.1.1 Theory
A real valued random variable $Z$ is (right-)heavy-tailed when $\operatorname{\mathbb{E}}[\exp(\lambda Z)]=\infty$ for all $\lambda>0$. Otherwise $Z$ is light-tailed.
Examples of heavy tailed distributions include Student’s $T$ and Pareto distributions. An example of a light tailed distribution is the Gaussian distribution. Jaini et al. (2020) prove the following result.
Theorem 2
@@
In further results Jaini et al. derive the gradient that $T$ requires to produce desired tail-properties, and generalise their results to the multivariate case. Finally they also prove that several popular classes of NFs use Lipschitz transformations (see also Liang et al., 2022).
@@ 1.1.2 Heavy Tailed Base Distributions
Firstly, in the setting of density estimation, Jaini et al. (2020) set $Z_{i}\sim t_{\nu}$: a standard (location zero, scale one) Student’s T distribution with $\nu$ degrees of freedom. The $Z_{i}$ components are independent and identically distributed. The $\nu$ parameter is learned jointly with the NF parameters. Optimising their objective function requires evaluating the Student’s $T$ log density and its derivative with respect to $\nu$, which is straightforward.
An extension is to allow tail anisotropy—tail behaviour varying across dimensions—by setting (independently) $Z_{i}\sim t_{\nu_{i}}$, so the degrees of freedom can differ with $i$. In Laszkiewicz et al. (2022), two such approaches are proposed for density estimation. Marginal Tail-Adaptive Flows (mTAF) first learn $\nu_{i}$ values using a standard estimator (a version of the Hill, 1975 estimator). Then the distribution of $Z$ is fixed while learning $T$. Another feature of mTAF is it allows Gaussian marginals by using Gaussian base distributions for appropriate $Z_{i}$ components, and then keeping Gaussian and non-Gaussian components separate throughout the flow. Generalised Tail-Adaptive Flows (gTAF) train the $\nu_{i}$s and $\theta$ jointly. Liang et al. (2022) propose a similar method to gTAF for the setting of variational inference: Anisotropic Tail-Adaptive Flows (ATAF).
@@ 1.1.3 Other Methods
Here we discuss work on alternatives to a Student’s T base distribution. In the setting of density estimation, Amiri et al. (2022) consider two alternative base distributions: Gaussian mixtures, and generalized Gaussian. For the former, they argue that mixture distributions can in theory model any smooth density given enough components, and also that they are more stable in optimisation than heavy tailed base distributions. The latter has tails which can be heavier than Gaussian, but are lighter than Student’s T.
McDonald et al. (2022) propose COMET (copula multivariate extreme) flows for the setting of density estimation. This involves a preliminary stage of estimating marginal cumulative distribution functions (cdfs) for each component of the data. The body of each marginal is approximated using kernel density estimation, and the tails are taken as GPDs, with shape parameters chosen using maximum likelihood. The COMET flow then uses a Gaussian base distribution, followed by a copula transformation. This is an arbitrary normalizing flow followed by an elementwise logistic transformation so that output is in $[0,1]^{d}$. Finally the inverse marginal cdfs are applied to each component. In Appendix B.2 we analyse the resulting tail behaviour and prove that the output cannot capture the full range of heavy tailed behaviour. (More formally, Theorem 11 proves that the output is not in the Fréchet domain of attraction, which is defined in Section 2.4.) Despite this our experiments in Section 4 find COMET flows often have good empirical performance, which we discuss in Section 5.
The extreme value theory literature also proposes methods to jointly model heavy tails and the body of a distribution. See Huser et al. (2024) (Section 4) for a review of several approaches. We comment on one particularly relevant method: Papastathopoulos and Tawn (2013); Naveau et al. (2016); de Carvalho et al. (2022) use a cdf $G\circ H(y)$ which composes the GPD cdf $H$ with a carrier function $G:[0,1]\to[0,1]$ which is taken to be a simple parametric function. Similarly, Stein (2021) composes several carrier functions. Our approach is similar, effectively using normalizing flows in place of carrier functions, and covering the multivariate case. Naveau et al. (2016) proves that certain conditions on the carrier function provide useful theoretical properties, and it would be interesting to explore analogous results for our method in future.
@@ 2.1 Normalizing Flows
Typically a parametric transformation $T(z;\theta)$ is used, and $q_{z}(z)$ is a fixed density, such as that of a $\mathcal{N}(0,I)$ distribution. However some previous work uses a parametric base density $q_{z}(z;\theta)$. (So $\theta$ denotes parameters defining both $T$ and $q_{z}$.) For instance several methods from Section 1.1 use a Student’s $T$ base distribution with variable degrees of freedom.
@@ 2.2 Density Estimation
This is equivalent to minimising a Monte Carlo approximation of the Kullback-Leibler divergence $KL[p(x)||q_{x}(x;\theta)]$. The objective gradient can be numerically evaluated using automatic differentiation. Thus optimisation is possible by stochastic gradient methods. This approach remains feasible under a Student’s $T$ base distribution as its log density and required derivatives are tractable.
@@ 2.3 Variational Inference
A generalisation of the above is to use a base distribution with parameters. In particular, ATAF (Liang et al., 2022) needs to learn degrees of freedom for Student’s $T$ distributions. The above approach remains feasible by using a Student’s $T$ sampling scheme which allows application of the reparameterization trick, as detailed in Appendix D.2.
For more background on this form of variational inference see Rezende and Mohamed (2015); Blei et al. (2017); Murphy (2023).
@@ 2.4 Extreme Value Theory
The Fréchet domain of attraction with shape parameter $\lambda$, $\Theta_{\lambda}$ is the set of distributions resulting in $\lambda>0$ under Pickands theorem.
@@ 3 Methods and Theory
We propose producing normalizing flow samples $R\circ T_{\text{body}}(z)$, where $z$ is a $\mathcal{N}(0,1)$ sample and $T_{\text{body}}$ is a standard normalizing flow transformation, and $R$ is defined shortly. If $T_{\text{body}}$ is a Lipschitz transformation, then the input to $R$ has light tails by Theorem 2. Thus the resulting architecture avoids the problem of passing extreme values as inputs to any neural network layers. The final transformation $R$ should be able to output heavy tails of any desired tail weight. This section presents our proposal for a suitable transformation $R$.
Section 3.1 describes our proposed transformation for the univariate case, and summarises our theoretical results on its tail behaviour (with full details in the appendices). Sections 3.23.4 give further details to produce a practical general-purpose method. Technical details are included in the appendices. This includes a discussion of alternative transformations and related results in the literature, in Appendix C.
@@ 3.1 TTF Transformation
We use the notation $s=\operatorname{sign}(z)$, with $\lambda_{s}=\lambda_{+}$ for $s=1$ and $\lambda_{s}=\lambda_{-}$ for $s=-1$. The transformation is based on $\operatorname{erfc}$, the complementary error function. This is a special function, reviewed in Appendix A.1, which is tractable for use in automatic differentiation using standard libraries. The parameters $\lambda_{+}>0,\lambda_{-}>0$ control tail weights for the upper and lower tails, allowing us to capture tail weight asymmetry in these two tails. The parameters $\mu\in\mathbb{R}$ and $\sigma>0$ are location and scale parameters. We found these helped performance, although similar effects could be achieved in principle by adjusting $T_{\text{body}}$.
To perform density evaluation via (2) we need to evaluate the inverse and derivative of (3). These, and some other properties, are provided in Appendix A.3.
Composing $R$ with existing normalizing flows permits the output to have heavy tails with parameterised weights. This is shown by the following argument. Most normalizing flows use a Gaussian base distribution and a Lipschitz transformation. By Theorem 2 the output has light tails. Since the NF transformation can be the identity, it is capable of producing Gaussian tails. A property of (3) is that if $X$ has Gaussian tails, then $R(X)$ is in the Fréchet domain of attraction (see Definition 3) with $\lambda_{+},\lambda_{-}$ controlling the tail shape parameters. A special case is that $\mathcal{N}(0,1)$ tails produce GPD output with shape parameters $\lambda_{+},\lambda_{-}$. These properties are proved in Appendix B.3 (Theorem 12).
@@ 3.2 Multivariate Transformation
Here we discuss extending our univariate transformation $R:\mathbb{R}\to\mathbb{R}$ to the multivariate case. In this paper we follow the simple approach of transforming each marginal with its own $\mu,\sigma,\lambda_{+},\lambda_{-}$ parameters. In some cases it can be known that a particular marginal is light tailed. Then we could simply perform an identity transformation instead. However, in practice we find that in such cases fixing $\lambda_{+},\lambda_{-}$ to low values (in particular $\lambda=1/1000$) suffices.
@@
3.3 Tail Parameter Initialisation
The initialisation of the tail parameters $\lambda_{+},\lambda_{-}$ is often important to the stability of the method. For density estimation in the presence of heavy tailed data, the parameters must be initialised sufficiently high, such that large observations are not mapped to vanishingly low probability regions of the base distribution. If this does occur, it is possible to get numerical overflow during optimisation. For variational inference a related problem can occur. If some tail parameters are too high, then we can sample points of very low target density and get numerical overflow. In practice, we find that initialising $\lambda$ uniformly from $[0.05,1]$ provides sufficient stability.
@@ 3.4 Two Stage Procedure
As discussed in Section 3.3, joint optimisation of $R$ and $T_{\text{body}}$ can require careful initialisation of the $\lambda$ parameters. As an alternative, we propose a two stage procedure for density estimation: TTF fix. Here the tail weight $\lambda_{-},\lambda_{+}$ parameters of $R$ are estimated in an initial step and then fixed while optimising $T_{\text{body}}$ (and the $\mu,\sigma$ parameters of $R$.) This can be viewed as first transforming the data using $R$ to remove heavy tails, and then fitting a standard normalizing flow to the transformed data. Similar approaches appear in McDonald et al. (2022) and Laszkiewicz et al. (2022).
Shape parameter estimators exist in the EVT literature, which we can apply to each marginal tail. We follow Laszkiewicz et al. (2022) in using the Hill double-bootstrap estimator (Danielsson et al., 2001; Qi, 2008). McDonald et al. (2022) perform maximum likelihood estimation on the highest and lowest 5% of data, to produce tail parameters for the positive and negative tails respectively.
We do not consider a similar two stage procedure for VI. This is because preliminary estimation of tail weights from the unnormalised target distribution is not straightforward. However recent work on static analysis of probabilistic programs (Liang et al., 2024) provides progress in this direction.

### v2

@@ Abstract
Normalizing flows are a flexible class of probability distributions, expressed as transformations of a simple base distribution. A limitation of standard normalizing flows is representing distributions with heavy tails, which arise in applications to both density estimation and variational inference. A popular current solution to this problem is to use a heavy tailed base distribution. We argue this can lead to poor performance due to the difficulty of optimising neural networks, such as normalizing flows, under heavy tailed input. We propose an alternative, “tail transform flow” (TTF), which uses a Gaussian base distribution and a final transformation layer which can produce heavy tails. Experimental results show this approach outperforms current methods, especially when the target distribution has large dimension or tail weight.
@@ 1 Introduction
Modelling distributions with heavy tails is critical in many applications such as climate (Zscheischler et al., 2018), contagious diseases (Cirillo and Taleb, 2020) and finance (Gilli and Këllezi, 2006). Indeed, in the context of density estimation, results from extreme value theory (EVT) (Coles, 2001; Embrechts et al., 2013) show distributional tails can often be modelled using a particular heavy tailed distribution, the generalized Pareto distribution (GPD). Heavy tailed targets can also arise naturally in Bayesian inference, requiring heavy tailed approximate distributions in variational inference (Liang et al., 2022) —and providing challenges for MCMC methods (see e.g. Yang et al., 2024).
Figure 1: Our method models extremes in the final layer, not the base distribution.
However standard NFs do not model heavy tails well. In particular, Jaini et al. (2020) prove that Gaussian tails cannot be mapped to GPD tails under Lipschitz transformations. (This is a special case of their main result, which we include as Theorem 2 below.) Many NFs use Lipschitz transformations of Gaussian base distributions, implying they produce distributions poorly approximating heavy tailed distributions.
Based on this result, several authors have proposed using heavy tailed base distributions. This includes the tail adaptive flow (TAF) methods of Laszkiewicz et al. (2022) for density estimation, and similar methods for variational inference (Liang et al., 2022). We argue that using heavy tailed base distributions has an important drawback. Typical NF transformations involve passing the input $z$ through a neural network. However neural network optimisation can perform poorly under heavy tailed input—as this can induce heavy tailed gradients, which are known to be problematic (Zhang et al., 2020). We verify this problem empirically in Section 4.1 for NFs (and in Appendix F for a simpler neural network regression example.)
To improve performance we propose an alternative: use a Gaussian base distribution with a final non-Lipschitz transformation in $T$. We refer to this as the tail transform flow (TTF) approach. The difference is illustrated in Figure 1. In Section 3.1 we provide a suitable final transformation (3), motivated by extreme value theory. This is easy to implement in automatic differentiation libraries as it’s based on a standard special function: the complementary error function. We also prove that this transformation converts Gaussian tails to heavy tails with tunable tail weights, and hence provides the capability of producing heavy tails when used with standard flows.
We perform experiments showing our TTF approach outperforms current methods, especially when the target distribution has large dimension or tail weight. We concentrate on density estimation, investigating both synthetic and real data, but also include a small-scale variational inference experiment.
@@
  • We illustrate the problem of using extreme inputs in NFs.
  • We introduce the TTF transformation—equation (3)—and prove it converts Gaussian tails to heavy tails with tunable tail weights. This provides the capability of producing heavy tails when used with standard NFs.
  • We provide practical methodology using TTF in a normalizing flow architecture.
  • We demonstrate improved empirical results for density estimation (synthetic and real data examples) and variational inference (an artificial target example) compared to standard NFs, and other NF methods for heavy tails.
@@ 1.1.1 Theory
A real-valued random variable $Z$ is (right-)heavy-tailed when $\operatorname{\mathbb{E}}[\exp(\lambda Z)]=\infty$ for all $\lambda>0$. Otherwise $Z$ is light-tailed.
Examples of heavy tailed distributions include Student’s $T$ and Pareto distributions. An example of a light tailed distribution is Gaussian distributions.
Under Definition 1 the following result holds.
Theorem 2 (Jaini et al., 2020)
@@
Jaini et al. (2020) also generalise this result to the multivariate case, and prove that several popular classes of NFs use Lipschitz transformations (see also Liang et al., 2022).
@@ 1.1.2 Heavy Tailed Base Distributions
Firstly, for density estimation, Jaini et al. (2020) set $Z_{i}\sim t_{\nu}$: a standard (location zero, scale one) Student’s T distribution with $\nu$ degrees of freedom. The $Z_{i}$ components are independent and identically distributed. The $\nu$ parameter is learned jointly with the NF parameters. Optimising their objective function requires evaluating the Student’s $T$ log density and its derivative with respect to $\nu$, which is straightforward.
An extension is to allow tail anisotropy—tail behaviour varying across dimensions—by setting (independently) $Z_{i}\sim t_{\nu_{i}}$, so the degrees of freedom can differ with $i$. In Laszkiewicz et al. (2022), two such approaches are proposed for density estimation. Marginal Tail-Adaptive Flows (mTAF) first learn $\nu_{i}$ values using a standard estimator (a version of the Hill, 1975 estimator). Then the distribution of $Z$ is fixed while learning $T$. Another feature of mTAF is it allows Gaussian marginals by using Gaussian base distributions for appropriate $Z_{i}$ components, and then keeping Gaussian and non-Gaussian components separate throughout the flow transformation. Generalised Tail-Adaptive Flows (gTAF) train the $\nu_{i}$s and $\theta$ jointly. Liang et al. (2022) propose a similar method to gTAF for the setting of variational inference: Anisotropic Tail-Adaptive Flows (ATAF).
@@ 1.1.3 Other Methods
Here we discuss work on alternatives to a Student’s T base distribution. In the setting of density estimation, Amiri et al. (2022) consider two alternative base distributions: Gaussian mixtures, and generalized Gaussian. Although the former are light tailed (see Appendix A), they argue that mixture distributions can in theory model any smooth density given enough components, and also that they are more stable in optimisation than heavy tailed base distributions. The latter has tails which can be heavier than Gaussian, but are lighter than Student’s T.
McDonald et al. (2022) propose COMET (copula multivariate extreme) flows for the setting of density estimation. This involves a preliminary stage of estimating marginal cumulative distribution functions (cdfs) for each component of the data. The body of each marginal is approximated using kernel density estimation, and the tails are taken as GPDs, with shape parameters chosen using maximum likelihood. The COMET flow then uses a Gaussian base distribution, followed by a copula transformation. This is an arbitrary normalizing flow followed by an elementwise logistic transformation so that output is in $[0,1]^{d}$. Finally the inverse marginal cdfs are applied to each component. In Appendix C.2 we analyse the resulting tail behaviour and prove that the output cannot capture the full range of heavy tailed behaviour. More formally, Theorem 11 proves that the output is not in the Fréchet domain of attraction—see Definition 3 below. This is in contrast to our proposed method, TTF, which we show in Appendix C.3 does produce in outputs in the Fréchet domain. Despite this our experiments in Section 4 find COMET flows often have good empirical performance, which we discuss in Section 5.
The extreme value theory literature also proposes methods to jointly model heavy tails and the body of a distribution. See Huser et al. (2025) (Section 4) for a review of several approaches. We comment on one particularly relevant univariate method: Papastathopoulos and Tawn (2013); Naveau et al. (2016); de Carvalho et al. (2022) use a cdf $G\circ H(y)$ which composes the GPD cdf $H$ with a carrier function $G:[0,1]\to[0,1]$ which is taken to be a simple parametric function. Similarly, Stein (2021) composes several carrier functions. Our approach is similar, effectively using NFs in place of carrier functions, and covering the multivariate case. Naveau et al. (2016) proves that certain conditions on the carrier function provide useful theoretical properties, and it would be interesting to explore analogous results for our method in future.
@@ 2.1 Normalizing Flows
Typically a parametric transformation $T(z;\theta)$ is used, and $q_{z}(z)$ is a fixed density, such as that of a $\mathcal{N}(0,I)$ distribution. However some previous work uses a parametric base density $q_{z}(z;\theta)$. (So $\theta$ denotes parameters defining both $T$ and $q_{z}$.) For instance some methods from Section 1.1 use a Student’s $T$ base distribution with variable degrees of freedom.
@@ 2.2 Density Estimation
This is equivalent to minimising a Monte Carlo approximation of the Kullback-Leibler divergence $KL[p(x)||q_{x}(x;\theta)]$. The objective gradient can be numerically evaluated using automatic differentiation. Thus optimisation is possible by stochastic gradient methods. This approach remains feasible under a Student’s $T$ base distribution since its log density and required derivatives are tractable.
@@ 2.3 Variational Inference
A generalisation of the above is to use a base distribution with parameters. In particular, ATAF (Liang et al., 2022) needs to learn degrees of freedom for Student’s $T$ distributions. The above approach remains feasible by using a Student’s $T$ sampling scheme which allows application of the reparameterization trick, as detailed in Appendix G.4.
For more background on this form of VI see Rezende and Mohamed (2015); Blei et al. (2017); Murphy (2023).
@@ 2.4 Extreme Value Theory
The Fréchet domain of attraction with shape parameter $\lambda$, $\Theta_{\lambda}$, is the set of distributions resulting in $\lambda>0$ under Pickands theorem.
@@ 3 Methods and Theory
We propose producing normalizing flow samples $R\circ T_{\text{body}}(z)$, where $z$ is a $\mathcal{N}(0,1)$ sample, $T_{\text{body}}$ is a standard normalizing flow transformation and $R$ is a final transformation. If $T_{\text{body}}$ is a Lipschitz transformation, then the input to $R$ has light tails by Theorem 2. Thus the resulting architecture avoids the problem outlined in Section 1 of passing extreme values as inputs to any neural network layers. The final transformation $R$ should be able to output heavy tails of any desired tail weight. This section presents our proposal for a suitable transformation $R$.
Section 3.1 describes our proposed transformation for the univariate case. Section 3.2 summarises our theoretical results on its tail behaviour. Sections 3.33.4 give further details to produce a practical general-purpose method, and Section 3.5 comments on universality. Proofs and technical details are given in the appendices. This also includes a discussion of alternative transformations and related results in the literature, in Appendix D.
@@ 3.1 TTF Transformation
We use the notation $s=\operatorname{sign}(z)$, with $\lambda_{s}=\lambda_{+}$ for $s=1$ and $\lambda_{s}=\lambda_{-}$ for $s=-1$. The transformation is based on $\operatorname{erfc}$, the complementary error function. This is a special function, reviewed in Appendix B.1, which is tractable for use in automatic differentiation using standard libraries. The parameters $\lambda_{+}>0,\lambda_{-}>0$ control tail weights for the upper and lower tails (also often referred to as right and left tails respectively). This allows us to model asymmetry in tail behaviour. The parameters $\mu\in\mathbb{R}$ and $\sigma>0$ are location and scale parameters. We found these helped performance, although similar effects could be achieved in principle by adjusting $T_{\text{body}}$.
To perform density evaluation via (2) we need to evaluate the inverse and derivative of (3). These, and some other properties, are provided in Appendix B.3.
3.2 Theory
Appendix C.3 proves results on the asymptotic properties of (3). Informally, if $X$ has Gaussian tails then $R(X)$ is in the Fréchet domain of attraction (see Definition 3) with $\lambda_{+},\lambda_{-}$ controlling the tail shape parameters. A special case is that $\mathcal{N}(0,1)$ tails produce GPD output with shape parameters $\lambda_{+},\lambda_{-}$.
So composing $R$ with existing NFs permits the output to have heavy tails with parameterised weights. This is shown by the following argument. Most NFs use a Gaussian base distribution and a Lipschitz transformation. By Theorem 2 the output has light tails. Since the NF transformation can be the identity, it is capable of producing Gaussian tails.
@@ 3.3 Multivariate Transformation
To extend our univariate transformation $R:\mathbb{R}\to\mathbb{R}$ to the multivariate case, we simply transform each marginal with its own $\mu,\sigma,\lambda_{+},\lambda_{-}$ parameters. In some cases we know particular marginals are light tailed. Then we could simply perform an identity transformation instead. However, we find fixing $\lambda_{+},\lambda_{-}$ to low values (we use $1/1000$) suffices.
@@ 3.4 Two Stage Procedure
Joint optimisation of $R$ and $T_{\text{body}}$ can require careful initialisation of the $\lambda$ parameters. (Details of how we do so are provided in Appendix G.3.) As an alternative, we propose a two stage procedure for density estimation: TTFfix. Here the tail weight $\lambda_{-},\lambda_{+}$ parameters of $R$ are estimated in an initial step and then fixed while optimising $T_{\text{body}}$ (and the $\mu,\sigma$ parameters of $R$.) This can be viewed as first transforming the data using $R$ to remove heavy tails, and then fitting a standard normalizing flow to the transformed data. Similar approaches appear in McDonald et al. (2022) and Laszkiewicz et al. (2022).
Shape parameter estimators exist in the EVT literature, which we can apply to each marginal tail. We follow Laszkiewicz et al. (2022) in using the Hill double-bootstrap estimator (Danielsson et al., 2001; Qi, 2008). Note that this enforces $\lambda_{-}=\lambda_{+}$. Alternatively McDonald et al. (2022) perform maximum likelihood estimation on the highest and lowest 5% of data, to produce tail parameters for the positive and negative tails respectively.
We do not consider a similar two stage procedure for VI, as preliminary estimation of tail weights from the unnormalised target distribution is not straightforward. However recent work on static analysis of probabilistic programs (Liang et al., 2024) provides progress in this direction.
3.5 Universality
It’s been proved that some NFs have a universality property: “the flow can learn any target density to any required precision given sufficient capacity and data” (Kobyzev et al., 2020). In Appendix E we show that many NF universality results are preserved when the TTF transformation is added as a final layer.
As we’ve already seen, the situation under bounded capacity is different. Standard NFs cannot produce heavy tailed distributions (Jaini et al. 2020, reviewed as Theorem 2 above). However adding our transformation does permit these (see Section 3.2).
So theoretically our method improves the set of distributions which can be modelled under bounded capacity without sacrificing expressiveness in the limit of infinite capacity. The next section shows this is reflected by improved empirical performance modelling heavy tailed data.